OpenStudy (anonymous):

3x - 18 = 2x - 20 answer please

OpenStudy (anonymous):

x+2 = 0 x= -2

OpenStudy (lilg132):

3x - 18 = 2x - 20 bring variables to one side and numbers on the other 3x - 2x = -20 + 18 x = -2

OpenStudy (anonymous):

thanks guys